Circular No. 8694

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2006 A4–A7, B2–B4, E2, F3 (SOHO) Additional comets have been found on SOHO website images (cf. IAUC

8679) — all being Kreutz sungrazers except for C/2006 B4 (Meyer group) and for C/2006 E2 and C/2006 F3 (Marsden group). C/2006 A4 is described by K. Battams as very faint and diffuse. C/2006 A5 had a bright coma that “evaporated” quickly in the C2 images, with a tail that “lingered for a long time”, being longer than 1^{o} on Jan. 5.612 UT at ∼ 4R_{⊙}; in C3 images, the comet was still brightening when at mag perhaps −1.5 on Jan. 5.112 at 11.6R_{⊙}, showing a bright, thin tail, but it was at mag ∼ 2.9 and fading fast by the time it reached 7R_{⊙}. The other comets were stellar in appearance. Peak magnitudes: C/2006 A6, 6.3, Jan. 10.431, 10.8R_{⊙}; C/2006 A7, 5.8, Jan. 12.738, 11.1R_{⊙}; C/2006 B3, 4.3, Jan. 22.612, 10.4R_{⊙}. C/2006 B2 and C/2006 F3 were very small and faint. C/2006 B4, at mag ≈ 7, was slightly elongated in the direction of motion. C/2006 E2 was extremely faint and slightly diffuse in C2 images (stellar in C3).

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 A4 Jan. 2.904 19^{h}04.^{m}2 −25^{o}26′ C3/2 RM 2006-F43
C/2006 A5  4.096 19 15.1 −27 53 C3/2 TH 2006-F43
C/2006 A6  10.071 19 40.1 −24 21 C3  TH 2006-F43
C/2006 A7  12.431 19 49.2 −24 12 C3  SF 2006-F43
C/2006 B2  20.321 20 22.8 −21 51 C3  HS 2006-F44
C/2006 B3  21.654 20 38.1 −23 16 C3  RK 2006-F44
C/2006 B4  26.496 20 37.2 −17 18 C2  HS 2006-F44
C/2006 E2 Mar. 15.096 23 50.5 − 1 26 C3/2 RM 2006-F50
C/2006 F3  26.101 0 23.1 + 4 13 C2  TH 2006-F50

V1647 ORIONIS V. Venkat and B. G. Anandarao, Astronomy and Astrophysics Divi-

sion, Physical Research Laboratory, Ahmedabad, report near-infrared photometry of V1647 Ori undertaken at Mt. Abu Observatory’s 1.2-m telescope (+ NICMOS infrared array) on Feb. 27, revealing magnitudes J = 14.6, H = 11.7, and K′ = 9.7. These magnitudes are close to the 2MASS values and fainter than their own values from 2005 Dec. 3 (Ojha et al. 2006,

MNRAS, in press; http://xxx.lanl.gov/abs/astro-ph/?0602044). It

appears therefore that, following its EXor-like eruption (IAUC 8289) that lasted more than 24 months (cf. IAUC 8600; 8681), the star may be near its pre-eruptive state.

2006 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green